A play ground is 1 yard longer than twice its width. John wants to increase the length and width buy 5 feet each to enlarge the area by 90 square feet. what will be the area of the new pen?
Answer by mananth(16949)

let width be x
length = 2x+1
...
increase by 5 ft
x+5 width
2x+1+5 = length
2x+6.
...
(x+5)(2x+6)-x(2x+1)=90
2x^2+6x+10x+30-2x^2-x=90
15x+30=90
-30
15x+30-30=90-30
15x=60
x=4 feet
New width = x+5 = 9 feet
New length = 2x+6 = 14 feet.
..